CASE 12 MID-ATLANTIC SPECIALTY, INC: Financial Risk 12/1/17
Copyright 2018 Foundation of the American College of Healthcare Executives. Not for sale.
Model without Questions, Student Version
Case 12 deals with basic financial risk concepts of a medical group practice facing alternative
investment opportunities.
The student version of the model is a format guide only because the objective is for students to attempt
basic calculations themselves. The key to student success in this case lies in correct calculations and
interpretation of the data.
Estimated one-year return distributions:
State of the 1-Year Healthcare Inverse Biotech S&P
Economy Probability T-Bill Fund ETF Fund 500
Poor 0.1 7% -8% 18% -19% -15%
Below average 0.2 7% 2% 23% 0% 0%
Average 0.4 7% 14% 7% 13% 15%
Above average 0.2 7% 25% -3% 31% 30%
Excellent 0.1 7% 33% 2% 50% 45%
Historical return distributions:
1-Year Healthcare Inverse Biotech S&P
T-Bill Fund ETF Fund 500
Year 1 7% -8% 18% -19% -15%
Year 2 7% 2% 23% 0% 0%
Year 3 7% 14% 7% 13% 15%
Year 4 7% 25% -3% 31% 30%
Year 5 7% 33% 2% 50% 45%
  Stand-alone return and risk:
1-Year Healthcare Inverse Biotech S&P
T-Bill Fund ETF Fund 500
Expected return [E(R)]
Variance (Var)
Standard deviation (SD)
Coefficient of variation (CV)
  Portfolio return and risk:
E(R) of E(R) of
Portfolio Portfolio
State of the Healthcare Fund / Healthcare Fund /
Economy Inverse ETF Biotech Fund
Poor
Below average
Average
Above average
Excellent
Portfolio Portfolio
Healthcare Fund / Healthcare Fund /
Inverse ETF Biotech Fund
Expected return [E(R)]
Variance (Var)
Standard deviation (SD)
Coefficient of variation (CV)
Correlation coefficient
  Market characteristic lines:
S&P 1-Year Healthcare Inverse Biotech
500 T-Bill Fund ETF Fund
Intercept
Market beta
Data point
Data point
Data point
Data point
Data point
Market Characteristic Lines
  Security market line:
1-Year Healthcare Inverse Biotech
T-Bill Fund ETF Fund
Risk-free rate
Required market return
Market beta
Investment required return
Investment expected return
Difference
Security Market Line
Answer #1

Historical returns for the various investment options are given below,

1-Year Healthcare Inverse Biotech S&P
T-Bill Fund ETF Fund 500
Year 1 7% -8% 18% -19% -15%
Year 2 7% 2% 23% 0% 0%
Year 3 7% 14% 7% 13% 15%
Year 4 7% 25% -3% 31% 30%
Year 5 7% 33% 2% 50% 45%

For the year 6 the expected return from the investment as per different economic scenario is given below

Estimated one-year return distributions:
State of the 1-Year Healthcare Inverse Biotech S&P
Economy Probability T-Bill Fund ETF Fund 500
Poor 0.1 7% -8% 18% -19% -15%
Below average 0.2 7% 2% 23% 0% 0%
Average 0.4 7% 14% 7% 13% 15%
Above average 0.2 7% 25% -3% 31% 30%
Excellent 0.1 7% 33% 2% 50% 45%

Expected return for each investment in year 6 will be,

1 Year T bill = .1 * 7 + .2 * 7 + .4 * 7 + .2 * 7 + .1 * 7 = .7 + 1.4 + 2.8 + 1.4 + .7 = 7 %

Health Care Fund = .1 * -8 + .2 * 2 + .4 * 14 + .2 * 25 + .1 * 33 = -.8 + .4 + 5.6 + 5 + 3.3 = 13.5 %

Inverse ETF = .1 * 18 + .2 * 23 + .4 * 7 + .2 * -3 + .1 * 2 = 1.8 + 4.6 + 2.8 -.6 + .2 = 8.8 %

Biotech Fund = .1 * - 19 + .2 * 0 + .4 * 13 + .2 * 31 + .1 * 50 = -1.9 + 0 + 5.2 + 6.2 + 5 = 14.5 %

S & P 500 = .1 * - 15 + .2 * 0 + .4 * 15 + .2 * 30 + .1 * 45 = -1.5 + 0 + 6 + 6 + 4,5 = 15 %

Variance of each of the investment can be calculatted as,

1 Year T bill = VAR(7,7, 7,7,7,7) = 0

Health Care Fund = VAR(8, 2,14 ,25,33,13.5) = 221.3

Inverse ETF = VAR(18 ,23, 7, -3 ,2 ,8.8) = 94.7

Biotech Fund = VAR(19, 0 ,13 ,31,50,14.5) = 573.2

S & P 500 = VAR(-15 ,0 ,15 ,30 ,45 ,15) = 450

Standard deviation for each of the investment,

1 Year T bill = STDEV(7,7, 7,7,7,7) = 0

Health Care Fund = STDEV(8, 2,14 ,25,33,13.5) = 14.8

Inverse ETF = STDEV(18 ,23, 7, -3 ,2 ,8.8) = 9.7

Biotech Fund = STDEV(19, 0 ,13 ,31,50,14.5) = 23.9

S & P 500 = STDEV(-15 ,0 ,15 ,30 ,45 ,15) = 21.2

Average return for each of the investment,

1 Year T bill = AVERAGE(7,7, 7,7,7,7) = 7

Health Care Fund = AVERAGE(8, 2,14 ,25,33,13.5) = 13.2

Inverse ETF =AVERAGE(18 ,23, 7, -3 ,2 ,8.8) = 9.3

Biotech Fund =AVERAGE(19, 0 ,13 ,31,50,14.5) = 14.9

S & P 500 = AVERAGE(-15 ,0 ,15 ,30 ,45 ,15) = 15

Coefficient of variation (CV) for each investment is the deviation per unit of return,

1 Year T bill = 0/7 = 0

Health Care Fund = 14.8/13.2 = 1.12

Inverse ETF = 9.7/9.3 = 1.04

Biotech Fund = 23.9/14.9 = 1,6

S & P 500 = 21.2/15 = 1.4

N.B - As part of HOMEWORKLIB RULES we can answer first four of any multipart question.